OpenStudy (anonymous):

what is the solution in this linear equation of this system 3x+y=-2 y=-x+4

OpenStudy (anonymous):

[i] 3x +y = -2 [ii] y = -x +4 rearranging [ii] by adding x to both sides x +y = 4 now solve the system by subtracting [i] from [ii] 3x +y -(x +y) = -2 -4 3x +y -x -y = -6 2x = -6 x = -3 solve for y in either [i] or [ii] y = -x +4 = -(-3) +4 = 3 +4 = 7 the solution of the system is (x, y) = (-3, 7),

OpenStudy (anonymous):

thank you for showing me the steps i thank that